OpenStudy (anonymous):

find power series (maclaurin series): integral [ e^(-x)^(2) ] dx

OpenStudy (adamaero):

well you know the series of e^x

OpenStudy (anonymous):

is that correct for e^x

OpenStudy (anonymous):

i meant e^-x

OpenStudy (zarkon):

are you looking for the series to \[\int e^{-x^2}dx\]

OpenStudy (anonymous):

affirmative.

OpenStudy (zarkon):

find the series to \[\Large e^{-x^2}\] then integrate term by term

OpenStudy (anonymous):

can you elaborate

OpenStudy (zarkon):

can you find the series of \(e^{-x^2}\)?

OpenStudy (anonymous):

no

OpenStudy (zarkon):

replace x in the expansion of \(e^x\) with \(-x^2\)

OpenStudy (anonymous):

okay...

OpenStudy (zarkon):

then integrate each term

OpenStudy (anonymous):

in this case i would prefer to only eat for a day

OpenStudy (zarkon):

\[\int e^{-x^2}dx=\int\sum_{n=0}^{\infty}\frac{(-x^2)^n}{n!}dx=\int\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{n!}dx=\sum_{n=0}^{\infty}\int\frac{(-1)^nx^{2n}}{n!}dx\]

OpenStudy (anonymous):

very illuminating, thank you very much!

OpenStudy (zarkon):

you still need to integrate...I'll leave that to you